OpenStudy (mathhelp346):

3 1/3 * 7 5/6 * 2 2/5 someone answer it please and show how you got it

OpenStudy (chaise):

3 1/3 x 7 5/6 x 2 2/5 You multiply the denominator by the number out the front, add the numerator and divide by the denominator. 10/3 x 47/6 x 12/5 Use a calculate to finish it up, and you get 62 2/3 or 188/3

OpenStudy (anonymous):

Assuming these are compound fractions: \[(3+1/3) * (7 + 5/6) * (2 + 2/5)\] \[= (10/3)*(47/6)*(12/5)\] \[= (10*47*12) / (3*6*5) \] \[= 5640 / 90\] \[= 188/3\]

OpenStudy (mathhelp346):

i got 21 13/15 is that right?

OpenStudy (mathhelp346):

eh

OpenStudy (mathhelp346):

ummm

OpenStudy (anonymous):

Unfortunately not if they are compound fractions, do you have any workings?

OpenStudy (mathhelp346):

uh..

OpenStudy (anonymous):

\[\left(3+\frac{1}{3}\right)*\left(7+\frac{5}{6}\right)*\left(2+\frac{2}{5}\right)=\frac{37}{5} \]

OpenStudy (mathhelp346):

Thank you!

OpenStudy (anonymous):

Thanks for the medal. Here is the solution in detail:\[\left(\frac{10}{3}\right)*\left(\frac{47}{6}\right)*\left(\frac{12}{5}\right)=\frac{10*47*12}{3*6*5}=\frac{5640}{90}=\frac{37}{5} \]

OpenStudy (anonymous):

Sorry, how did you get from \[\frac{5640}{90}\]to \[\frac{37}{5}\]

OpenStudy (mathhelp346):

What did you divide to get from 5640 to 188?

OpenStudy (anonymous):

Sorry. Take the medal away. The answer with further verification is the same as Mathsboy's:\[\frac{5640}{90}=\frac{188}{3} \]

OpenStudy (anonymous):

No need to take the medal away, the working was decent, I was just a little confused.

OpenStudy (anonymous):

Divide top and bottom by 30.

OpenStudy (mathhelp346):

It's impossible to take the medal away anyways...right

OpenStudy (mathhelp346):

Ok thanks you two

OpenStudy (anonymous):

I used Mathematica to formulate the answer. This program uses double == sign as the normal = sign is used in everyday mathematics. So the answer is formed by a series of copy and pastes following the deletion of one of the "=" symbols. Mathematica never makes a mistake if the input is correct.

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

I've made that mistake too many times in Mathematica.